A construction engineer wanted to conduct research study and wishes to estimate, with 90% confidence, the proportion of people living in metro manila who own a home built by a constructor. A previous study showed that 35% of those interviewed had their houses built by a construction firm. The engineer wishes to be accurate within 2.5% of the true proportion. Find the minimum sample size necessary.
Answer #1

Solution :

Given that,

\hat p = 35% = 0.35

1 - \hat p = 1 - 0.35 = 0.65

margin of error = E = 2.5% = 0.025

At 90% confidence level the z is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.645 / 0.025)2 * 0.35 * 0.65

= 985

sample size = 985